OpenStudy (anonymous):

I dont understand at all(*@&$#)@ The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (mokeira):

when two lines are parallel they have the same slope....you know that right?

OpenStudy (anonymous):

Yeah of course

OpenStudy (mokeira):

equation of a line is \[y=mx+c\] \[y=\frac{ 1 }{2 }x+6\] so m=1/2 now we can form the equation of the parallel line \[m=\frac{ \Delta y }{ \Delta x}=\frac{ y_1-y_2 }{x_1-x_2 }\] so, \[\frac{ 1 }{2 }=\frac{ y--2 }{ x-0 }\] cross multiply

OpenStudy (anonymous):

If those are the answers it would be A right? y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (texaschic101):

y = 1/2x + 6 slope here is 1/2. Parallel lines have same slope. now use y = mx + b slope(m) = 1/2 (0,-2)...x = 0 and y = -2 now we sub, and since we already know the slope(m), we need to find the y intercept(b) -2 = 1/2(0) + b -2 = b so your parallel equation is : y = 1/2x - 2

OpenStudy (anonymous):

How did I miss that, Thank you so much guys

OpenStudy (texaschic101):

no problem .....lol...if I had $1 for every mistake I made, I would be rich :)

OpenStudy (mokeira):

\[\frac{ 1 }{2 }=\frac{ y+2 }{ x }\] cross multiply \[x=2(y+2)\] \[x=2y+4\] \[2y=x-4\] divide both sides by 2 \[y=\frac{ 1 }{ 2 }x -2\]

OpenStudy (mokeira):

lol... me too @texaschic101